# Byzantine Resilient Federated Multi-Task Representation Learning

### v2

@@ IV-A Experimental Setup
Datasets. We evaluated our algorithm on two real-world datasets, CIFAR-10 and Federated Extended MNIST (FEMNIST). CIFAR-10 consists of 60,000 $32\times 32$ color images in 10 classes, with 50,000 training images and 10,000 test images. FEMNIST consists of 145,600 $28\times 28$ pixel images of handwritten characters, with 124,800 training images and 20,800 test images. It incorporates both uppercase and lowercase letters, and the data is partitioned based on the individual writers of characters, offering a broader classification task.
Model. We use the AlexNet architecture, which is composed of five convolutional blocks followed by a linear layer for the shared representation and an additional linear layer for the personalized client heads. Each convolutional block includes a convolutional layer with a stride and padding of $1$, a ReLU activation function, and a max-pooling layer with a kernel size of $2$. We optimize the model using stochastic gradient descent (SGD) with a learning rate $\eta=0.01$ and momentum $\beta=0.9$.
Hyperparameters. We initialized all models with random parameters. For CIFAR-10, training was conducted for $T=100$ communication rounds, while FEMNIST models were trained for $T=200$ rounds. In each round, $20\%$ of clients ($\alpha=0.2$) were selected to perform local updates and receive the aggregated model from the server. Each selected client executed 10 local epochs of SGD with momentum for head training and one epoch for updating the global representation.
Data was distributed in a non-iid manner, with each client assigned different classes. We considered $100$ and $1000$ clients for CIFAR-10 and $150$ clients for FEMNIST. Among them, $20$ out of $100$, $100$ out of $1000$, and $50$ out of $150$ clients were designated as malicious. By varying the proportion of Byzantine clients, we analyzed their impact on learned model.
Benchmark Algorithms. We evaluated our algorithm with benchmark approaches, including FedAvg [1], FedPer [17], and FedRep [3]. FedPer and FedRep are similar approaches that learn a shared representation with personalized heads, but FedPer updates both simultaneously, maintaining an equal number of updates per local round. In contrast, FedAvg learns a single global model without personalization. As a best-case benchmark, we considered these methods without malicious clients. Additionally, we analyzed two variants: (i) introducing malicious clients (Byzantine FedPer/FedRep/FedAvg), (ii) incorporating geometric median-based aggregation (FedAvg+GM, FedPer+GM), and (iii) incorporating Krum-based aggregation (FedAvg+Krum, FedPer+Krum). For meta-test, we consider an additional baseline, Naive, which trains new clients independently without leveraging the MTRL framework.
Performance metrics. We evaluate the performance of our algorithm by calculating the average classification accuracy across all clients in the presence of malicious clients.
@@ IV-C Experiment Results
(a) CIFAR-10
(b) FEMNIST
(c) Meta test on CIFAR-10
(d) Meta test on FEMNIST
@@
Figure 2:   FedRep,   Byzantine FedRep,   BR-MTRL+GM (proposed),   BR-MTRL+Krum (proposed)   Naive. Average test accuracy across all clients: Figure 2(a) shows the results for CIFAR-10 with 100 clients with 20 malicious clients. Figure 2(b) presents the results for FEMNIST with 150 clients with 20 malicious clients. We compared our proposed BR-MTRL with the FedRep algorithm under benign conditions (i.e., no malicious client) and FedRep subjected to Byzantine attacks (Byzantine FedRep). Meta test for new clients: Figures 2(c) and 2(d) present the average test accuracy across 10 new clients (CIFAR-10 and FEMNIST) after fine-tuning the local heads for 10 epochs using the shared model. We first learned the shared model using source clients and then used the shared model to optimize the local head of new clients.
CIFAR-10 FEMNIST
(# Clients n, # Classes per Client S, # Byzantine Clients) (100,2,20) (100,5,20) (1000,2,100) (150,3,50)
FedAvg [1] 40.08 62.62 36.66 50.42
FedPer[17] 90.81 82.29 79.23 76.82
FedRep [3] 87.70 80.17 80.79 76.63
Byzantine FedAvg 13.44 16.45 22.33 17.80
Byzantine FedPer 52.72 31.11 57.26 35.21
Byzantine FedRep 52.08 32.69 58.17 35.25
FedAvg+GM 36.38 41.77 31.84 38.43
FedAvg+Krum 31.52 42.25 30.08 38.20
FedPer+GM 72.37 50.08 63.03 52.47
FedPer+Krum 71.90 61.21 68.17 53.12
BR-MTRL+GM 72.94 58.55 75.92 55.28
BR-MTRL + Krum 80.25 61.80 72.76 54.07
@@
TABLE II: Average test accuracy across all clients with some clients under byzantine attacks for different algorithms.
Test accuracy. As shown in Figures 2(a) and 2(b) and Table II, the FedRep and FedPer algorithms consistently achieved higher accuracy than FedAvg across various experiments, irrespective of the number of clients or classes per client, demonstrating the effectiveness of personalized FL. However, the performance of these three baselines (without any malicious clients) significantly declined under Byzantine attacks. This decline validates that (personalized) FL frameworks susceptible to Byzantine attacks will lead to unreliable models with poor performance. Increasing the number of classes per client from two to five resulted in a substantial accuracy drop of $21.61\%$. This is because Byzantine clients can negatively impact the global model’s performance across a greater number of classes. Furthermore, reducing the proportion of Byzantine clients relative to the total number of clients improves performance, as seen in the comparison between experiments with $(100,2,20)$ and $(1000,2,100)$, i.e., $20\%,10\%$ malicious clients, respectively, where the latter exhibits greater resilience.
Our experimental results further demonstrate that the proposed approaches (BR-MTRL+GM and BR-MTRL+Krum) achieves a higher accuracy than both the GM and Krum-adapted versions of FedPer. This is because FedPer updates the server model at the same frequency as client updates, making it more susceptible to Byzantine clients, which can adversely affect learning accuracy. In contrast, our approach updates the shared model only once after updating the client heads, leading to improved overall performance. All experimental results consistently showcase the improved performance of proposed approach. Krum outperforms GM in experiments with $100$ clients; however, has a lower accuracy when the number of clients increases. Krum requires to know the number of malicious clients and computational complexity is higher.
@@
Results for Meta-test. We tested the transferability of the proposed approach. We first learn the shared model $\phi$ using the source clients and then utilize this model to train the local heads of new, previously unseen clients. For each new client, we fine-tune the client head for $10$ epochs, using the transferred shared model. For CIFAR-10, we randomly selected 10 clients as new clients and trained the shared model using the remaining 90 clients, including 20 Byzantine clients. Similarly, for FEMNIST, we selected 10 new clients and trained the shared model with the remaining 140 clients, of which 10 were Byzantine. We evaluated the average accuracy for four approaches. To highlight the benefits of MTRL for transfer learning, we considered a Naive setting, where each client trains their model independently, resulting in poor accuracy. Figures 2(c) and 2(d) illustrate the results: FedRep, without Byzantine attacks, demonstrates effective transferability of the shared model to new clients. In contrast, Byzantine MTRL, with malicious clients, exhibits significantly lower performance, highlighting the vulnerability of FL systems. Finally, BR-MTRL, our proposed approach, improves learning accuracy by employing GM and Krum aggregation, validating its effectiveness.
@@ V Conclusion
In this paper, we proposed a Byzantine-resilient federated MTRL framework to learn robust shared representations across related yet different tasks. Our approach employs an alternating gradient descent strategy, where the shared model is updated while keeping the local head fixed, and vice versa. To ensure Byzantine resilience, we incorporated geometric median and Krum-based aggregation at the server, mitigating the impact of malicious clients sending corrupted updates. We implemented our federated learning approach on AWS, simulating a real-world federated setting. Extensive experiments on CIFAR-10 and FEMNIST validate the effectiveness of our approach and its transferability to new clients with limited data.

### v3

@@ IV-A Experimental Setup
Datasets. We evaluated our algorithm on two real-world datasets, CIFAR-10 and FEMNIST. CIFAR-10 consists of 60,000 $32\times 32$ color images in 10 classes, with 50,000 training images and 10,000 test images. FEMNIST consists of 145,600 $28\times 28$ pixel images of 26 handwritten characters, with 124,800 training images and 20,800 test images.
Model. We use a deep neural network, which is composed of two convolutional blocks followed by a linear layer for the shared representation and an additional linear layer for the personalized client heads. Each convolutional block includes a convolutional layer with a stride and padding of $1$, a ReLU activation function, and a max-pooling layer with a kernel size of $2$. We optimize the model using stochastic gradient descent (SGD) with a learning rate $\eta=0.01$ and momentum $\beta=0.9$.
Hyperparameters. We initialized all models with random parameters. For CIFAR-10, training was conducted for $T=100$ communication rounds, while FEMNIST models were trained for $T=200$ rounds. In each round, all the clients perform local updates and receive the aggregated model from the server. Each client executed 10 local epochs of SGD with momentum for head training and one epoch for updating the global representation.
The data distribution followed a highly non-i.i.d., pathological partitioning scheme, where each client was assigned a fixed subset of classes and, for each class, randomly sampled 750 images. We considered $100$ and $1000$ clients for CIFAR-10 and $150$ clients for FEMNIST. Among them, $20$ out of $100$, $100$ out of $1000$, and $50$ out of $150$ clients were designated as malicious. By varying the proportion of Byzantine clients, we analyzed their impact on the learned model.
Benchmark Algorithms. We evaluated our algorithm with benchmark approaches, including FedAvg [1], FedPer [14], and FedRep [5]. FedPer and FedRep are similar approaches that learn a shared representation with personalized heads; FedPer updates both $\phi,h$ simultaneously within each round $t$ while FedRep updates $\phi,h$ in an alternating fashion. In contrast, FedAvg learns a single global model without personalization. As a best-case benchmark, we considered these methods without malicious clients. Additionally, we analyzed these variants: (i) introducing malicious clients by adding noise to local representation (SR FedPer/FedRep/FedAvg), (ii) introducing malicious clients by misclassifying the image label (ML FedPer/FedRep/FedAvg), (iii) incorporating GM aggregation (FedAvg+GM, FedPer+GM), and (iv) incorporating Krum aggregation (FedAvg+Krum, FedPer+Krum). For meta-test, we consider an additional baseline, Naive, which trains new clients independently without leveraging the MTRL framework.
Performance metrics. We evaluate the performance of our algorithm by calculating the average classification accuracy across all benign clients in the presence of malicious clients.
@@ IV-C Experiment Results
(a) CIFAR-10 (SR)
(b) FEMNIST (SR)
(c) Meta test: CIFAR-10 (SR)
(d) Meta test: FEMNIST (SR)
(e) CIFAR-10 (ML)
(f) FEMNIST (ML)
(g) Meta test: CIFAR-10 (ML)
(h) Meta test: FEMNIST (ML)
@@
Figure 2:   FedRep,   Byzantine FedRep,   BR-MTRL+GM (proposed),   BR-MTRL+Krum (proposed)   Naive. Average test accuracy across all clients: Figure 2(a) and 2(e) show the results for CIFAR-10 with 100 clients with 20 malicious clients under SR and ML attacks respectively. Figure 2(b) and  2(f) present the results for FEMNIST with 150 clients with 50 malicious clients under SR and ML attacks respectively. We compared our proposed BR-MTRL with the FedRep algorithm under benign conditions (i.e., no malicious client) and FedRep subjected to Byzantine attacks (Byzantine FedRep). Meta test for new clients: Figures 2(c), 2(d), 2(h) and 2(g) present the average test accuracy across 10 new clients (CIFAR-10 and EMNIST) after fine-tuning the local heads for 10 epochs using the shared model. We first learned the shared model using source clients and then used the shared model to optimize the local head of new clients.
Method CIFAR-10 FEMNIST
(100,2,20) (100,5,20) (1000,2,100) (150,3,50)
FedAvg [1] 38.56 55.39 60.81 61.46
FedPer [14] 82.03 75.59 75.10 93.77
FedRep [5] 86.78 82.11 81.11 94.69
SR FedAvg 10.43 15.52 11.71 5.78
SR FedPer 52.44 38.51 54.00 37.37
SR FedRep 54.88 35.41 54.26 43.47
SR FedAvg+Krum 15.78 39.30 32.28 11.50
SR FedAvg+GM 38.11 55.18 35.45 11.45
SR FedPer+Krum 74.72 67.48 70.05 83.11
SR FedPer+GM 76.26 70.12 73.16 85.73
SR BR-MTRL+Krum 76.28 74.21 74.68 80.28
SR BR-MTRL+GM 81.48 82.03 75.36 88.47
ML FedAvg 27.90 28.25 13.86 32.65
ML FedPer 28.07 46.66 48.83 60.95
ML FedRep 57.54 58.04 60.31 42.72
ML FedAvg+Krum 31.23 36.02 45.88 48.16
ML FedAvg+GM 30.86 51.67 52.40 54.91
ML FedPer+Krum 71.65 63.15 68.80 72.00
ML FedPer+GM 74.85 69.53 68.94 73.41
ML BR-MTRL+Krum 73.96 70.01 74.25 80.05
ML BR-MTRL+GM 79.97 76.15 80.05 89.11
@@
TABLE II: Average test accuracy (%) across all benign clients under Scaled Random (SR) and Mislabeling (ML) Byzantine attacks for different algorithms. $(n,S,b)$ denote number of clients, classes per client, and number of Byzantine clients.
Test accuracy. As shown in Figures 2(a), 2(b), 2(e), 2(f) and Table II, the FedRep and FedPer algorithms consistently achieved higher accuracy than FedAvg for all experiments, demonstrating the effectiveness of personalized FL. However, the performance of these three baselines (without malicious clients) significantly declined under Byzantine attacks. This decline validates that (personalized) FL frameworks susceptible to Byzantine attacks will lead to unreliable models with poor performance. The SR setting induces a more significant decline, especially as the number of classes per client increases from $(100,2,20)$ to $(100,5,20)$, indicating that greater task heterogeneity amplifies the propagation of corrupted updates. In contrast, the ML setting exhibits an increase, suggesting that personalized representations partially mitigate label-based corruption. Additionally, scalability enhances robustness across both scenarios: when the proportion of malicious clients decreases from $20\%$ $(100,2,20)$ to $10\%$ $(1000,2,100)$, the aggregation process benefits from the dominance of benign contributions. Our experimental results further demonstrate that the proposed approaches, BR-MTRL+GM, achieve a higher accuracy than both the GM and Krum-adapted versions of FedPer. This is because the proposed alternating gradient descent approach within each round is more effective than FedPer’s simultaneous updates of the client head $h$ and server model $\phi$. All experimental results consistently showcase the improved performance of the proposed approach over the two attacks. Furthermore, GM outperforms Krum in most of the experiments. One potential reason is that Krum’s output is overly conservative, as it is exactly one of the client updates, discarding information from other honest clients. The GM-based solution is more effective, especially when the number of malicious clients is less. Further, Krum requires knowing the number of malicious clients, and has higher computation.
@@
Results for Meta-test. We tested the transferability of the proposed approach. We first learn the shared model $\phi$ using the source clients and then utilize this model to train the local heads of new, previously unseen clients. For each new client, we fine-tune the client head for $10$ epochs, using the transferred shared model. For CIFAR-10, we randomly selected 10 clients as new clients and trained the shared model using the remaining 90 clients, including 20 Byzantine clients. Similarly, for FEMNIST, we selected 10 new clients and trained the shared model with the remaining 140 clients, of which 50 were Byzantine. We evaluated the average accuracy for four approaches. To highlight the benefits of MTRL for transfer learning, we considered a Naive setting, where each client trains their model independently, resulting in poor accuracy. Figures 2(c), 2(d), 2(h) and 2(g) illustrate the results: FedRep, without Byzantine attacks, demonstrates effective transferability of the shared model to new clients. In contrast, Byzantine FedRep, with malicious clients with two type of attacks, exhibits significantly lower performance, highlighting the vulnerability of FL systems. Finally, BR-MTRL, our proposed approach, improves learning accuracy by employing GM and Krum aggregation, validating its effectiveness u under Byzantine attacks.
@@ V Conclusion
In this paper, we proposed a Byzantine-resilient federated MTRL framework to learn robust shared representations across related yet different tasks. Our approach employs an alternating gradient descent strategy, where the shared model is updated while keeping the local head fixed, and vice versa. To ensure Byzantine resilience, we incorporated geometric median and Krum-based aggregation at the server, mitigating the impact of malicious clients sending corrupted updates. We implemented our federated learning approach on AWS, simulating a real-world-like federated setting. Numerical experiments on CIFAR-10 and FEMNIST validate the effectiveness of our approach and its transferability to new clients with limited data.